# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v1

@@ Abstract
Large language models (LLMs), trained on diverse data effectively acquire a breadth of information across various domains. However, their computational complexity, cost, and lack of transparency hinder their direct application for specialised tasks. In fields such as clinical research, acquiring expert annotations or prior knowledge about predictive models is often costly and time-consuming. This study proposes using LLMs to elicit expert prior distributions for predictive models. This approach also provides an alternative to in-context learning, where language models are tasked with making predictions directly. We compare LLM-elicited and uninformative priors, evaluate whether LLMs truthfully generate parameter distributions, and propose a model selection strategy for in-context learning and prior elicitation. Our findings show that LLM-elicited prior parameter distributions significantly reduce predictive error compared to uninformative priors in low-data settings. Applied to clinical problems, this translates to fewer required biological samples, lowering cost and resources. Prior elicitation also consistently outperforms and proves more reliable than in-context learning at a lower cost, making it a preferred alternative in our setting. We demonstrate the utility of this method across various use cases, including clinical applications. For infection prediction, using LLM-elicited priors reduced the number of required labels to achieve the same accuracy as an uninformative prior by 55%, at 200 days earlier in the study.
@@ 1 Introduction
In fields such as clinical research, labelling data can be costly and time-consuming [1]. For example, in healthcare monitoring, patients’ data is recorded through wearables or remote monitoring devices and observations are continuously collected. However, the labelling of such data for clinical outcomes, such as diagnosing illnesses or collecting and studying samples, can be challenging. Further, estimating uncertainty on clinical predictions is paramount [2], especially when available training data is sparse.
Expert knowledge about a problem domain can provide prior distributions on model parameters. This can improve a model’s performance when trained with few data points and enable estimates of uncertainties on parameters and predictions. However, expert knowledge about model parameter distributions is often hard to obtain due to the challenges of performing prior elicitation from experts [3, 4].
Due to their diverse training data, language models offer an opportunity to begin resolving these issues. By considering a language model as an “expert” in a given field, we can extract insights about a predictive task that might lead to improved predictive models. This, and the abilities of language models to reason about problem settings outside of their intended use cases [5, 6, 7, 8, 9, 10], have inspired their use in performing prior elicitation [11, 12, 13]. Specifically, Selby et al. [12] show, by considering prior predictive distributions, that GPT-4 [10] can provide parameter distributions that allow for zero-shot performance (making predictions without training examples) of Bayesian models that compare to models trained on tens of data points. Further, in Requeima et al. [13], the authors prompt language models with text and collected data points to elicit posterior predictive distributions for regression tasks, which they find outperforms Gaussian processes with radial basis function kernels on some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations and a given text description that contains prior information from a human expert. Given the rate of development of language models [14, 15, 16, 17], it is conceivable that their knowledge improves and broadens with time, motivating the construction of a probabilistic framework to reason about elicited prior task knowledge from language models, as well as understand their use for predictive tasks.
@@
In-context learning allows for the direct use of language models for predictive tasks. Here, language models are provided task descriptions and demonstration examples from a training data set (in text form) and are prompted to make predictions on unseen data [18, 19]. In this sense, in-context learning is similar to eliciting a prior distribution on a predictive model’s parameters and calculating its posterior predictive distribution, where the main differences arise from: (1) Language models are notoriously weak at performing arithmetic on numerical values [20]; (2) Language models lack transparency and may not be learning from demonstrations in a predictable or Bayesian way [21]; (3) Language models may not be able to faithfully provide their internal knowledge of a task when specifying a prior parameterisation for a separate predictive model or use that same prior knowledge in-context. To study the trade-offs between using in-context learning and prior elicitation we aim to use Bayes factor for model selection, and propose an approach to estimate a language model’s internal predictive prior and posterior models.
We are motivated by the following questions, and make the following contributions:
@@
    Finding: We find these distributions differ significantly from those that we elicit (Section 3.4).
@@
    Method: By additionally providing the language model with training demonstrations, we additionally extract the LLM’s in-context predictive parameter posterior distribution (Section 2.3.2). We can then perform Monte Carlo sampling on the extracted prior and measure its difference to the extracted posterior.
    Finding: We find that the language is inconsistent in its approximation of Bayesian inference, and that its reliability depends on the task (Section 3.5).
@@
    Question: When is the performance better with in-context learning?
@@
For implementing our methods and reproducing our results, we provide details about the code [^1] https://github.com/alexcapstick/llm-elicited-priors in Appendix A.1 and complete experimental information in Appendix A.5.
@@ 2.1 Defining the task descriptions
Our goal in this section is to define the space of prompts that we use to elicit priors and gain predictions from a language model. We start by defining the space of all tasks as $\mathcal{T}$. An individual task, such as a specific classification or regression problem, will be referred to as $T\in\mathcal{T}$. Since a description of a task can be rephrased many times without changing its meaning, we need to consider a space of possible task descriptions, referred to as $\mathcal{I}$, for each task. We say that the probability of a random text sequence $I\in\mathcal{I}$ being a relevant task description for task $T\in\mathcal{T}$ is: $p(I|T)$.
We assume that all relevant task descriptions have equal probability, i.e. if $I_{i}$ and $I_{j}$ fully describe $T$, then $p(I_{i}|T)=p(I_{j}|T)$. We also assume that irrelevant task descriptions have $p(I|T)=0$. This means that for a task $T$, all of the mass of $p(\mathcal{I}|T)$ is uniformly spread across the relevant task descriptions only.
In reality, the number of relevant task descriptions would be too large to cover since each word in the task description could be replaced with a synonym without losing the meaning of the full description. In Appendix A.11, we show that there are diminishing returns as you increase the number of ways of describing a task. For our experiments, we use $100$ relevant task descriptions to cover this space.
In practice, to produce $100$ task descriptions, we employ a language model to rephrase a single human written system role and user role $10$ times each and take their product. This means that using $100$ task descriptions is not particularly time-consuming and could be increased significantly by providing many base descriptions or prompting the language model to produce many more variants.
@@ 2.2.1 Background
Bayesian inference allows us to incorporate our knowledge about a predictive task before observing data. This is done through the use of the prior knowledge $p(\theta)$ (over a predictive model parameterised by $\theta$), which can be our own intuition about which features are more important for a prediction and in which way they correlate. Additionally, using a prior allows us to report uncertainty over new predictions $(\tilde{x},\tilde{y})$, through the posterior predictive distribution $p(y|\tilde{x},D)$. If the prior is well-specified, using a prior can also reduce the number of data points required to construct a comparably accurate predictive model.
A key problem in machine learning is defining this prior distribution $p(\theta)$ given expert knowledge. Previous work studied how this distribution can be elicited from experts [3], which can be challenging and costly. In parallel, research has shown that large language models (LLMs) encode knowledge about a surprising variety of tasks [5, 6, 7, 8, 9, 10], which could be used to generate information about a predictive task [11, 12, 13], possibly supplemented with scientific discussions [22].
Choi et al. [11] explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset [23]. This suggests the language model’s utility in providing a prior distribution over the parameters of a predictive model, inspiring us to explore the possibility of using LLMs for eliciting prior parameter distributions and allowing a direct comparison to in-context learning.
@@
Additionally, Zhu and Griffiths [24] describe using iterated learning to elicit prior distributions over language models for causal learning, proportion estimation, and predicting everyday quantities. The authors employ a Markov chain Monte Carlo method to access an LLM’s prior knowledge by performing successive inference using previous model outputs that tend towards a limiting distribution. By doing this, the authors assume that in-context learning is a form of Bayesian inference. However, in their work Zhu and Griffiths [24] do not consider how prior elicitation from LLMs can be used for predictive models or test whether an LLM is performing Bayesian inference, both of which we study in this work.
@@ 2.2.2 Our contribution: LLM-elicited priors
In our experiments, our mixture is over $100$ components, using several different task descriptions. Mixtures of Gaussians are able to approximate any probability density with enough components, allowing us to approximate the language model’s belief about the task, even if non-Gaussian. We would also expect the language model to have learnt how to parameterise a Gaussian distribution from its training, making them good candidates. However, mixtures of other distributions can be used instead and may be more useful for particular tasks.
@@ 2.3.1 Background
In the above notation, it is assumed that the output from the language model contains only the predicted label $\tilde{y}$. Alternatively, the log-likelihood of the tokens representing the possible labels can be measured from the language model’s output, leading to a modified notation as in Dong et al. [19].
@@
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use [18, 19, 25, 26, 15]. In Brown et al. [18], the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. [27] study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. [21] argue that in-context learning allows the language model to observe the label space, distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. The majority of the existing works focus on natural language tasks [19], leaving numerical problems less explored and a focus of our work.
@@
In Hollmann et al. [28], the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Additionally, Akyürek et al. [29] and Von Oswald et al. [30] discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. [29] prove that by construction, transformers are able to perform gradient descent and closed-form ridge regression. Akyürek et al. [29] also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. [30] show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal and Misra [31] propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
@@
To fully compare in-context learning and prior elicitation, we must consider two main questions: (1) Does the language model use the same prior in-context that we elicit? (2) Does in-context learning approximate Bayesian inference? To answer these, we aim to extract, in the case of a linear model, the in-context predictive model’s prior and posterior parameter distributions.
@@ 2.3.2 Our contribution: Extracting the in-context prior and posterior
By specifying the structure of the internal predictive model to be used by the language model $M$ within our task descriptions $I$ for a task $T$, we can write the internal model as: $f_{\varphi}:X\rightarrow Y$ where $f$ represents the class of the predictive model, $\varphi$ is the random variable associated with its parameters, and $X$ and $Y$ correspond to the predictive model’s domain and range respectively. Depending on the model class $f$ corresponds to, $y\in Y$ could be a probability distribution over a set of possible targets or the prediction of the target itself.
@@
Given a language model $M$ and task $T$, we construct the sample distribution (following Murphy [32]) as:
@@
Where $\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ refers to the maximum likelihood estimate of the parameters $\varphi$ given input feature $X$ and in-context predictions $f_{\varphi}(X)$ for a task description $I$. This is a sample from the estimated prior distribution over the parameters $\varphi$.
We can additionally provide $\mathcal{D}$ examples with the task description to get the posterior samples:
@@
These are distributions estimating $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. For a more detailed breakdown of the process, see Appendix A.2. In our experiments, to sample each $\varphi_{\text{MLE}}$, we choose to sample $X\sim\mathcal{U}[-5,5]$ five times for each of the $100$ task descriptions. Since we normalise our real data, these samples of $X$ will cover the domain of $f_{\varphi}$ with a high probability.
In the linear case (for both linear and binary logistic regression), this process is simple since we can calculate the maximum likelihood estimate of $y=\varphi_{0}+\varphi\cdot x$ in closed form as we have direct access to $p_{M,T}(y|x,\varphi)$, which is either a regression label or the logit of a class.

### v2

@@ Abstract
Large language models (LLMs), trained on diverse data effectively acquire a breadth of information across various domains. However, their computational complexity, cost, and lack of transparency hinder their direct application for specialised tasks. In fields such as clinical research, acquiring expert annotations or prior knowledge about predictive models is often costly and time-consuming. This study proposes the use of LLMs to elicit expert prior distributions for predictive models. This approach also provides an alternative to in-context learning, where language models are tasked with making predictions directly. In this work, we compare LLM-elicited and uninformative priors, evaluate whether LLMs truthfully generate parameter distributions, and propose a model selection strategy for in-context learning and prior elicitation. Our findings show that LLM-elicited prior parameter distributions significantly reduce predictive error compared to uninformative priors in low-data settings. Applied to clinical problems, this translates to fewer required biological samples, lowering cost and resources. Prior elicitation also consistently outperforms and proves more reliable than in-context learning at a lower cost, making it a preferred alternative in our setting. We demonstrate the utility of this method across various use cases, including clinical applications. For infection prediction, using LLM-elicited priors reduced the number of required labels to achieve the same accuracy as an uninformative prior by 55%, 200 days earlier in the study.
@@ 1 Introduction
In fields such as clinical research, labelling data can be costly and time-consuming [1]. For example, in healthcare monitoring, patient data is recorded through wearables or remote monitoring devices, and observations are continuously collected. However, labelling such data for clinical outcomes, such as diagnosing illnesses or collecting and studying samples, can be challenging. In addition, estimating the uncertainty in clinical predictions is paramount [2], especially when available training data is sparse.
Expert knowledge of a problem domain can provide prior distributions of model parameters. This can improve a model’s performance when trained with few data points and enable estimates of uncertainties on parameters and predictions. However, expert knowledge about model parameter distributions is often difficult to obtain due to the challenges of performing prior elicitation from experts [3, 4], especially in low-resource settings.
Due to their diverse training data, language models offer an opportunity to begin resolving these issues. By considering a language model as an “expert” in a given field, we can extract insights about a predictive task that might lead to improved predictive models. This, and the abilities of language models to reason about problem settings outside of their intended use cases [5, 6, 7, 8, 9, 10], have inspired their use in performing prior elicitation [11, 12, 13]. Specifically, Selby et al. [12] show, by considering prior predictive distributions, that GPT-4 [10] can provide parameter distributions that allow for zero-shot performance (making predictions without training examples) of Bayesian models that compare to models trained on tens of data points. Furthermore, in Requeima et al. [13], the authors prompt language models with text and collected data points to elicit posterior predictive distributions for regression tasks, which they find outperforms Gaussian processes with radial basis function kernels on some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations, and a given text description that contains prior information from a human expert. Given the rate of development of language models [14, 15, 16, 17], it is conceivable that their knowledge improves and broadens with time. This motivates the construction of a probabilistic framework to extract prior task knowledge from language models and evaluate the use of their elicited distributions for predictive tasks.
@@
In-context learning allows for the direct use of language models for predictive tasks. Here, language models are provided with task descriptions and demonstration examples from a training data set (in text form) and are asked to make predictions on unseen data [18, 19]. In this sense, in-context learning is similar to eliciting a prior distribution on a predictive model’s parameters and calculating its posterior predictive distribution, where the main differences arise from: (1) Language models are notoriously weak at performing arithmetic on numerical values [20]; (2) Language models lack transparency and may not be learning from demonstrations in a predictable or Bayesian way [21]; (3) Language models may not be able to faithfully provide their internal knowledge of a task when specifying a prior parameterisation for a separate predictive model or use that same prior knowledge in-context. (4) Language models incur a significant cost, and so their use for predictions in a high-frequency or low-resource setting is infeasible. To study the trade-offs between using in-context learning and prior elicitation, we aim to use the Bayes factor for model selection, and propose an approach to estimate a language model’s internal predictive prior and posterior models.
We are motivated by the following questions and make the following contributions:
@@
    Finding: We find that these distributions differ significantly from those elicited (Section 3.4).
@@
    Method: By additionally providing the language model with training demonstrations, we extract the posterior distribution of the LLM’s in-context parameters (Section 2.3.2). We can then perform Monte Carlo sampling on the extracted prior and measure its difference to the extracted posterior.
    Finding: We find that the language model is inconsistent in its approximation of Bayesian inference and that its reliability depends on the task (Section 3.5).
@@
    Question: When is performance better with in-context learning?
@@
To implement our methods and reproduce our results, we provide details on the code [^1] https://github.com/alexcapstick/llm-elicited-priors in Appendix A.1 and complete experimental information in Appendix A.5.
@@ 2.1 Defining the task descriptions
The goal of this section is to define the space of prompts that we use to elicit priors and extract in-context predictions from a language model. We start by defining the space of all tasks as $\mathcal{T}$. An individual task, such as a specific classification or regression problem, will be referred to as $T\in\mathcal{T}$. Since a description of a task can be rephrased many times without changing its meaning, we must consider the space of possible task descriptions, referred to as $\mathcal{I}$. We say that the probability of a random text sequence $I\in\mathcal{I}$ being a relevant task description for task $T\in\mathcal{T}$ is: $p(I|T)$.
We assume that all relevant task descriptions have equal probability, that is, if $I_{i}$ and $I_{j}$ fully describe $T$, then $p(I_{i}|T)=p(I_{j}|T)$. We also assume that irrelevant task descriptions have $p(I|T)=0$. This means that for a task $T$, all of the mass of $p(\mathcal{I}|T)$ is uniformly spread across the relevant task descriptions only.
In reality, the number of relevant task descriptions would be too large to cover, since each word in the task description could be replaced with a synonym without losing the meaning of the full description. In Appendix A.11, we show that there are diminishing returns as you increase the number of ways of describing a task. For our experiments, we use $100$ relevant task descriptions to cover this space.
In practice, to produce $100$ task descriptions, we employ a language model to rephrase a single human written system role and user role $10$ times each and take their product. This means that using $100$ task descriptions is not particularly time-consuming and could be significantly increased by providing many base descriptions or prompting the language model to produce many more variants.
@@ 2.2.1 Background
Bayesian inference allows us to incorporate our knowledge about a predictive task before observing the data. This is done using the prior knowledge $p(\theta)$ (over a predictive model parameterised by $\theta$), which can be our own intuition about which features are more important for a prediction and in what way they correlate. Additionally, using a prior allows us to report uncertainty over new predictions $(\tilde{x},\tilde{y})$, through the posterior predictive distribution $p(y|\tilde{x},D)$. If the prior is well-specified, using a prior can also reduce the number of data points required to construct a comparably accurate predictive model.
A key problem in machine learning is to define this prior distribution $p(\theta)$ given expert knowledge. Previous work studied how this distribution can be elicited from experts [3], which can be challenging and costly, or infeasible in low-resource settings. In parallel, research has shown that large language models (LLMs) encode knowledge about a surprising variety of domains [5, 6, 7, 8, 9, 10], which could be used to generate information about a predictive task [11, 12, 13], possibly supplemented with scientific discussions between experts and the language model[22].
Choi et al. [11] explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset [23]. This inspires us to explore the possibility of using LLMs for eliciting prior parameter distributions over linear models.
@@
Furthermore, Zhu and Griffiths [24] describe the use of iterated learning to elicit prior distributions from language models for causal learning, proportion estimation, and predicting everyday quantities. The authors employ a Markov chain Monte Carlo method to access an LLM’s prior knowledge by performing successive inference using previous model outputs that tend towards a limiting distribution. In doing this, the authors assume that in-context learning is a form of Bayesian inference. However, in their work Zhu and Griffiths [24] do not consider how prior elicitation from LLMs can be used for predictive models or test whether an LLM is performing Bayesian inference, both of which we study in this work.
@@ 2.2.2 Our contribution: LLM-elicited priors
In our experiments, our mixture has over $100$ components, using several different task descriptions. Mixtures of Gaussians are able to approximate any probability density with enough components, allowing us to approximate the language model’s belief about the task, even if non-Gaussian. We would also expect the language model to have learnt how to parameterise a Gaussian distribution from its training, making them good candidates. However, mixtures of other distributions can be used instead and may be more useful for particular tasks.
@@ 2.3.1 Background
In the above notation, it is assumed that the output of the language model contains only the predicted label $\tilde{y}$. Alternatively, the log-likelihood of the tokens representing the possible labels can be measured from the language model’s output, leading to a modified notation as in Dong et al. [19].
@@
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use [18, 19, 25, 26, 15]. In Brown et al. [18], the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. [27] study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. [21] argue that in-context learning allows the language model to observe the label space, the distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. Most of the existing work focuses on natural language tasks [19], leaving numerical problems less explored and a focus of our work.
@@
In Hollmann et al. [28], the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Furthermore, Akyürek et al. [29] and Von Oswald et al. [30] discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. [29] show that, by construction, transformers are capable of performing gradient descent and closed-form ridge regression. Akyürek et al. [29] also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. [30] show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal and Misra [31] propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
@@
To fully compare in-context learning and prior elicitation, we must consider two main questions: (1) Does the language model use the same prior in-context that we elicit? (2) Does in-context learning approximate Bayesian inference? To answer these questions, we aim to extract, in the case of a linear model, the in-context predictive model’s prior and posterior parameter distributions.
@@ 2.3.2 Our contribution: Extracting the in-context prior and posterior
By specifying the structure of the internal predictive model to be used by the language model $M$ within our task descriptions $I$ for a task $T$, we can write the internal model as: $f_{\varphi}:X\rightarrow Y$ where $f$ represents the class of the predictive model, $\varphi$ is the random variable associated with its parameters, and $X$ and $Y$ correspond to the domain and range of the predictive model, respectively. Depending on the model class $f$ corresponds to, $y\in Y$ could be a probability distribution over a set of possible targets or the prediction of the target itself.
@@
Given a language model $M$ and task $T$, we construct the sample distribution (following Murphy [32]) as follows:
@@
Where $\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ refers to the maximum likelihood estimate of the parameters $\varphi$ given input feature $X$ and in-context predictions $f_{\varphi}(X)$ for a task description $I$. This is a sample of the estimated prior distribution over the parameters $\varphi$.
We can also provide $\mathcal{D}$ examples with the task description to get posterior samples:
@@
These are distributions that estimate $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. For a more detailed breakdown of the process, see Appendix A.2. In our experiments, to sample each $\varphi_{\text{MLE}}$, we choose to sample $X\sim\mathcal{U}[-5,5]$ five times for each of the $100$ task descriptions. Since we normalise our real data, these samples of $X$ will cover the domain of $f_{\varphi}$ with a high probability.
In the linear case (for both linear and binary logistic regression), this process is simple since we can calculate the maximum likelihood estimate of $y=\varphi_{0}+\varphi\cdot x$ in closed form since we have direct access to $p_{M,T}(y|x,\varphi)$, which is either a regression label or the logit of a class.